OpenStudy (anonymous):

Find the integral from 0 to 1/4 of 1/square root of 1-4x^2 dx

hartnn (hartnn):

did you try the substitution, 2x = sin y ?

hartnn (hartnn):

then what would be 1-4x^2 = ... ?

OpenStudy (anonymous):

idk

hartnn (hartnn):

have you solved an integration question before ?

OpenStudy (anonymous):

yes, I'm trying to figure out how to get rid of the 4

hartnn (hartnn):

right, thats was the reason to substitute "2"x = sin y to get rid of that 4 once you find what \(\sqrt {1-4x^2}\) is, you'll come to know

OpenStudy (anonymous):

That is a known integral. \[ \int \frac{1}{\sqrt{1-4 x^2}} \, dx=\frac{1}{2} \sin ^{-1}(2 x) \]

OpenStudy (anonymous):

\[ \frac{1}{2} \sin ^{-1}(2 (1/4))- \frac{1}{2} \sin ^{-1}(0)=\frac{1}{2} \sin ^{-1}(1/2)=\frac 1 2 \frac \pi 6=\frac \pi {12} \]

OpenStudy (anonymous):

We use the fact that \[ \frac{d \sin ^{-1}(2 x)}{ dx}=\frac{2}{\sqrt{1-4 x^2}} \]